1.     Assume that velocity is constant, money growth is 5 percent per year, and real GDP is growing at 2 percent per year. 

Y=10,000 

T=2,500 

G=3,000 

C=1000+0.8(Y-T

I=2000-2000

Given the above information, compute

a.     , and S (i.e., private, government, and national saving),

b.     r  (the real interest rate),

c.     (the inflation rate),

d.     i  (the nominal interest rate).


1
Expert's answer
2021-07-27T20:20:01-0400

(a)

"S=Y-C"

"=10,000-[1000+0.8(10,000-2500)]"

"=3000"

(b)

"r=\frac{i-g}{1+g}"

"r= \frac {(2000-2000r)-3000}{1+3000}"

"r=-0.2%" %

(c)

"1+i=(1+r)(1+\pi)"

"1+2400=(1-0.2)(1+\pi)"

"\pi =2400"

(d)

"i=2000-2000r"

"=2000+400"

"=2400"